# A Testbed for Tidal Migration: the 3D Architecture of an Eccentric Hot Jupiter HD 118203 b Accompanied by a Possibly Aligned Outer Giant Planet

### v3

@@ Abstract
Characterizing outer companions to hot Jupiters plays a crucial role in deciphering their origins. We present the discovery of a long-period giant planet, HD 118203 c ($m_{c}=11.79^{+0.69}_{-0.63}\ \mathrm{M_{J}}$, $a_{c}=6.23^{+0.16}_{-0.15}$ AU) exterior to a close-in eccentric hot Jupiter HD 118203 b ($P_{b}=6.135\ \mathrm{days}$, $m_{b}=2.14\pm{0.12}\ \mathrm{M_{J}}$, $r_{b}=1.14\pm{0.029}\ \mathrm{R_{J}}$, $e_{b}=0.31\pm{0.007}$) based on twenty-year radial velocities. Using Rossiter-McLaughlin (RM) observations from the Keck Planet Finder (KPF), we measured a low sky-projected spin-orbit angle $\lambda_{b}=-11^{\circ}.7^{+7.6}_{-10.0}$ for HD 118203 b and detected stellar oscillations in the host star, confirming its evolved status. Combining the RM observation with the stellar inclination measurement, we constrained the true spin-orbit angle of HD 118203 b as $\Psi_{b}<33^{\circ}.5\ (2\sigma)$, indicating the orbit normal of the hot Jupiter nearly aligned with the stellar spin axis. Furthermore, by combining radial velocities and Hipparcos-Gaia astrometric acceleration, we constrained the line-of-sight mutual inclination between the hot Jupiter and the outer planet to be $9^{\circ}.8^{+16.2}_{-9.3}$ at $2\sigma$ level. HD 118203 is one of first hot Jupiter systems where both the true spin-orbit angle of the hot Jupiter and the mutual inclination between inner and outer planets have been determined. Our results are consistent with a system-wide alignment, with low mutual inclinations between the outer giant planet, the inner hot Jupiter, and the host star. This alignment, along with the moderate eccentricity of HD 118203 c, implies that the system may have undergone coplanar high-eccentricity tidal migration. Under this framework, our dynamical analysis suggests an initial semi-major axis of 0.3 to 3.2 AU for the proto-hot Jupiter.
@@ 1 Introduction
The spin-orbit angles (also known as stellar obliquities) of stars hosting hot Jupiters are considered a crucial indicator of their formation channel (Winn et al., 2010a; Albrecht et al., 2012; Dawson & Johnson, 2018; Rice et al., 2022). Spin-orbit angle is the angle between the star’s spin axis and the planet’s orbital axis. In scenarios involving tidal migration, where dynamic events cause eccentric orbits, the angular momentum of the proto-hot Jupiters would be altered, leading to high spin-orbit angles.
Additionally, stellar companions might cause misalignment between protoplanetary disks and their host stars (Batygin, 2013). Thus, planets could end up misaligned relative to their host stars even if they migrate through disk-driven migration. However, observations have suggested that the spin-orbit angle distribution may not only reflect the hot Jupiters’ origins but also sculpting by tidal realignment. Schlaufman (2010) and Winn et al. (2010a) indicate a significant dichotomy in the spin-orbit angles of stars hosting hot Jupiters based on the star’s temperature. They discovered that cool stars hosting hot Jupiters generally show low spin-orbit angles, likely because effective tidal realignment processes erase their initial misalignment. In contrast, hotter stars exhibit a broader range of spin-orbit angles, indicating that they may retain the initial spin-orbit angles from when hot Jupiters first migrated to their close-in orbit. Furthermore, Rice et al. (2022) found that the spin-orbit angles of eccentric hot Jupiters do not exhibit such a trend with stellar temperature, which is consistent with the high-eccentricity migration scenario from their models.
@@
HD 118203 (TOI-1271) is a sub-giant star ($M_{*}=1.25\pm{0.06}\ M_{\odot}$, $R_{*}=2.10\pm{0.06}\ R_{\odot}$) hosting a transiting hot Jupiter with an orbital period of 6.135 days, a mass of $2.166^{+0.074}_{-0.079}\ M_{J}$ and an eccentricity of 0.3 (da Silva et al., 2006; Pepper et al., 2020). In this paper, we first combined Rossiter-McLaughlin observations from the Keck Planet Finder (KPF) and stellar rotation periods derived from Transiting Exoplanet Survey Satellite (TESS) light curves to determine the true spin-orbit angle of HD 118203 b. Furthermore, we present the discovery of HD 118203 c, a long-period giant planet orbiting outside HD 118203 b. We characterized the three-dimensional orbital parameters of this outer planet by combining RVs and Hipparcos and Gaia astrometric data. HD 118203 is one of the first systems where the spin-orbit angle for both the hot Jupiter and the outer giant planet, and the mutual inclination between the hot Jupiter and outer planet have been determined. Recently, Maciejewski et al. (2024) presented an independent detection of HD 118203 c using long-baseline radial velocity data from HARPS-N (Cosentino et al., 2012) and the High-Resolution Spectrograph (HRS; Tull 1998) at the Hobby-Eberly Telescope (Ramsey et al., 1998). We compare our results with their findings in Section 5.3.
@@ 3.1 Literature values
HD 118203 is a sub-giant G0V star at an age of $5.32^{+0.96}_{-0.73}$ Gyr and a distance of $91.811\pm{0.236}$ pc (Pepper et al., 2020). The star has a mass of $1.25\pm{0.06}\ M_{\odot}$, a radius of $2.10\pm{0.06}\ R_{\odot}$, and a metallicity [Fe/H] of $0.223\pm{0.076}$ dex. Its spectroscopically determined projected rotation velocity, $v\sin i_{*}$, is $5.32\pm{0.5}$ km s-1 (Luck, 2017). Pepper et al. (2020) reported a rotation period of $20\pm{5}$ days from TESS light curves of sector 15 and 16. Recently, Castro-González et al. (2024) reported a period pf 6.1 days from TESS PDCSAP light curves, which matches with the orbital period of HD 118203 b. They argue it could be resulted from the magnetic interaction between the hot Jupiter and the host star.
@@ 5.2 RV-only Keplerian Fit
We performed the Markov Chain Monte Carlo (MCMC) exploration with emcee (Foreman-Mackey et al., 2013) to estimate parameter credible levels. Our MCMC analysis used 50 walkers and ran for $1.5\times 10^{6}$ steps per walker, achieving a maximum Gelman-Rubin(GR) statistic of 1.004. Figure 6 shows the best fit Keplerian solution. We derived a minimum mass for HD 118203 c from RV amplitudes as $m_{c}\sin{c}=11.7^{+0.87}_{-0.83}\ \rm M_{J}$, semi-major axis as $a_{c}=6.32^{+0.17}_{-0.18}$ AU and an eccentricity of $e_{c}=0.268^{+0.028}_{-0.026}$. The derived planetary parameters are given in Table 5.
@@ 5.3 Three-dimensional orbit fitting
Semi-major axis $a_{c}$ (AU) $6.23^{+0.16}_{-0.15}$ no prior
@@
Figure 8 presents the joint posterior distributions for the orbital inclination, mass and orbital period of HD 118203 c (see Appendix A.4 for additional parameters). Our results show that HD 118203 c is a giant planet with a mass of $m_{c}=11.79^{+0.69}_{-0.63}\ \mathrm{M_{J}}$ semi-major axis of $a_{c}=6.23^{+0.16}_{-0.15}$ AU and eccentricity $e_{c}=0.26^{+0.03}_{-0.02}$. We obtained the orbital inclination of the outer giant planet as $I_{c}=93^{\circ}.7^{+11.9}_{-13.8}$, suggesting that HD 118203 c orbits the star in an almost edge-on orbit. Our results are consistent with the values reported from Maciejewski et al. (2024) within $1\sigma$.
Figure 9: Panel a: sketch showing the orbital geometry for the mutual inclination between HD 118203 b&\&&c. The observer is looking down from the z-axis. The x-y plane is the plane of the sky, and the star is centered on the origin. The orbital inclination of a planet is defined as the angle between the planet orbital axis and line of sight direction (z-axis). The mutual inclination is the angle between the orbital axes of the two planets. Panel b: For a transiting planet with orbital inclination $I_{1}=90^{\circ}$ without knowing the longitude of acceding node, its orbital axis could fall anywhere along the sky plane (orange). For a planet with an orbital inclination $I_{2}\neq 90^{\circ}$ without knowing the longitude of acceding node, its orbital axis falls on the surface of a cone (blue) with a vertex angle equals to the orbital inclination (toward observer, $I_{2}<90^{\circ}$) or $\pi-$ orbital inclination (away from observer, $I_{2}>90^{\circ}$). The dots $P_{1}$ and $P^{\prime}_{1}$ represent two orbits that exhibit pro-grade and retrograde degeneracy: they both lie within the same plane but orbit in opposite directions.
Figure 10: (a) Orbital inclination of HD 118203 c ($I_{c}$, blue) , compared with that of HD 118203 b ($I_{b}$, red) and the line-of-sight stellar inclination ($i_{*}$, orange). Zero degrees correspond to the line of sight, same as z-axis in Figure 9. The stellar inclination ranges from $0^{\circ}$ to $180^{\circ}$ to accounts for scenarios where the stellar axis may point either toward ($<90^{\circ}$) or away ($>90^{\circ}$) from the observer. (b): Distribution of HD 118203 b&\&&c mutual inclination projected in line of sight direction. (c): Distribution of spin-orbit angle of HD 118203 c projected in line of sight direction.
@@ 5.4 Constraints on the planet-planet mutual inclination and spin-orbit angle for HD 118203 c
where $I_{i}$ and $\Omega_{i}$ is the inclination and longitude of ascending node for planet $i$. Typically, using transit and radial velocity data, we can infer the inclination of a transiting planet, but not its longitude of the ascending node. In such scenarios, the orbital axis of a transiting planet with an inclination close to $90^{\circ}$ might be oriented in any direction within the sky plane. Similarly, for a non-transiting planet with an inclination away from $90^{\circ}$, its orbital axis could lie on the surface of a cone, which is either towards ($I_{2}<90^{\circ}$) or away from ($I_{2}>90^{\circ}$) the observer. Furthermore, when two orbits have an inclination sum equal to $180^{\circ}$, as seen in $P_{1}$ and $P^{\prime}_{1}$ in Figure 9b, their orbital axes align along the same line but in opposite directions. This indicates that both orbits occupy the same plane but the orbit directions are opposite. This phenomenon is known as the retrograde-prograde degeneracy. Therefore, if we only have information about the orbital inclinations of two planets and the longitude of ascending nodes for one or both planets is unknown, we cannot determine the true mutual inclinations. Instead, we can only derive the mutual inclinations as projected along the line of sight.
@@ 6.2.1 Eccentricity excitation
However, our findings that the outer planet has an eccentricity of $e_{c}=0.26$, imply that we need to take into account the effects of the non-circular outer orbit. Naoz et al. (2011) explored a more generalized mechanism in which the outer orbit can be eccentric, called the eccentric Kozai-Lidov (EKL) mechanism. They found that considering an eccentric outer orbit requires an expansion of the perturbing function to octupole order, which is proportional to $(a_{1}/a_{2})^{3}$. In the octupole level of approximation, EKL mechanism can drive the eccentricity of the inner orbit to extremely high levels, and can even induce a complete reversal in the orientation of the inner orbit, flipping it from prograde to retrograde. The EKL mechanism is also applicable across a broader range of parameters, including scenarios with low initial mutual inclinations between the two objects. Li et al. (2014) demonstrated that when the two planets have eccentric orbits, the inner planet’s eccentricity can grow to large values even when the orbits are initially close to being coplanar. Petrovich (2015) investigated the consequence of this eccentricity excitation in the context of tides and found that subsequent high-eccentricity migration kept the two planets nearly coplanar. According to their study, hot Jupiters formed via this coplanar high-eccentricity migration (CHEM) are likely to have distant ( $a\gtrsim 5\rm{AU}$) and massive (three times more massive than the hot Jupiter) companions with relatively low mutual inclinations ( $\leq 20^{\circ}$) and moderate eccentricities (0.2-0.5). Our measurements for HD 118203 c, with properties of $a_{c}=6.23^{+0.16}_{-0.15}\ \rm{AU}$, $m_{c}=11.79^{+0.69}_{-0.63}\ M_{J}$, and $e_{c}=0.26^{+0.03}_{-0.02}$, align well with these predictions. Furthermore, our analysis indicates that $\approx 95\%$ values from the posterior distribution of the line-of-sight $\Delta I_{bc}$ are below $25^{\circ}$, consistent with the outer giant planet having a low mutual inclination relative to the inner hot Jupiter.
@@ 7 Conclusions
  • HD 118203 c is a long-period giant planet ($m_{c}=11.79^{+0.69}_{-0.63}\ \mathrm{M_{J}}$, $a_{c}=6.23^{+0.16}_{-0.15}$ AU, $e_{c}=0.26^{+0.03}_{-0.02}$) with moderate eccentricity outside the hot Jupiter. We constrained the orbital inclination of HD 118203 c ($I_{c}=93^{\circ}.7^{+11.9}_{-13.8}$ ) by combing RV and Hipparcos-Gaia astrometric accelerations. The line-of-sight mutual inclination between HD 118203 b &\&& c is $9^{\circ}.8^{+8.6}_{-6.7}$ at $1\sigma$ and $9^{\circ}.8^{+16.2}_{-9.3}$ at $2\sigma$ confidence level. The line-of-sight spin-orbit angle for HD 118203 c is $14^{\circ}.7^{+13.3}_{-10.2}$ at the $1\sigma$ confidence level and $14^{\circ}.7^{+26.4}_{-13.9}$ at the $2\sigma$ level confidence level. Our measurements demonstrate that the orbit of HD 118203 c is most likely aligned with both the host star spin axis and the orbit of the inner hot Jupiter.

### v4

@@ Abstract
Characterizing outer companions to hot Jupiters plays a crucial role in deciphering their origins. We present the discovery of a long-period giant planet, HD 118203 c ($m_{c}=11.79^{+0.69}_{-0.63}\ \mathrm{M_{J}}$, $a_{c}=6.28^{+0.10}_{-0.11}$ AU) exterior to a close-in eccentric hot Jupiter HD 118203 b ($P_{b}=6.135\ \mathrm{days}$, $m_{b}=2.14\pm{0.12}\ \mathrm{M_{J}}$, $r_{b}=1.14\pm{0.029}\ \mathrm{R_{J}}$, $e_{b}=0.31\pm{0.007}$) based on twenty-year radial velocities. Using Rossiter-McLaughlin (RM) observations from the Keck Planet Finder (KPF), we measured a low sky-projected spin-orbit angle $\lambda_{b}=-11^{\circ}.7^{+7.6}_{-10.0}$ for HD 118203 b and detected stellar oscillations in the host star, confirming its evolved status. Combining the RM observation with the stellar inclination measurement, we constrained the true spin-orbit angle of HD 118203 b as $\Psi_{b}<33^{\circ}.5\ (2\sigma)$, indicating the orbit normal of the hot Jupiter nearly aligned with the stellar spin axis. Furthermore, by combining radial velocities and Hipparcos-Gaia astrometric acceleration, we constrained the line-of-sight mutual inclination between the hot Jupiter and the outer planet to be $9^{\circ}.8^{+16.2}_{-9.3}$ at $2\sigma$ level. HD 118203 is one of first hot Jupiter systems where both the true spin-orbit angle of the hot Jupiter and the mutual inclination between inner and outer planets have been determined. Our results are consistent with a system-wide alignment, with low mutual inclinations between the outer giant planet, the inner hot Jupiter, and the host star. This alignment, along with the moderate eccentricity of HD 118203 c, implies that the system may have undergone coplanar high-eccentricity tidal migration. Under this framework, our dynamical analysis suggests an initial semi-major axis of 0.3 to 3.2 AU for the proto-hot Jupiter.
@@ 1 Introduction
The spin-orbit angles (also known as stellar obliquities) of stars hosting hot Jupiters are considered a crucial indicator of their formation channel (Winn et al., 2010a; Albrecht et al., 2012; Dawson & Johnson, 2018; Rice et al., 2022). Spin-orbit angle is the angle between the star’s spin axis and the planet’s orbital axis. In scenarios involving tidal migration, where dynamic events cause eccentric orbits, the angular momentum of the proto-hot Jupiters would be altered, leading to high spin-orbit angles. Additionally, stellar companions might cause misalignment between protoplanetary disks and their host stars (Batygin, 2013). Thus, planets could end up misaligned relative to their host stars even if they migrate through disk-driven migration. However, observations have suggested that the spin-orbit angle distribution may not only reflect the hot Jupiters’ origins but also sculpting by tidal realignment. Schlaufman (2010) and Winn et al. (2010a) indicate a significant dichotomy in the spin-orbit angles of stars hosting hot Jupiters based on the star’s temperature. They discovered that cool stars hosting hot Jupiters generally show low spin-orbit angles, likely because effective tidal realignment processes erase their initial misalignment. In contrast, hotter stars exhibit a broader range of spin-orbit angles, indicating that they may retain the initial spin-orbit angles from when hot Jupiters first migrated to their close-in orbit. Furthermore, Rice et al. (2022) found that the spin-orbit angles of eccentric hot Jupiters do not exhibit such a trend with stellar temperature, which is consistent with the high-eccentricity migration scenario from their models.
@@
HD 118203 (TOI-1271) is a sub-giant star ($M_{*}=1.25\pm{0.06}\ M_{\odot}$, $R_{*}=2.10\pm{0.06}\ R_{\odot}$) hosting a transiting hot Jupiter with an orbital period of 6.135 days, a mass of $2.166^{+0.074}_{-0.079}\ M_{J}$ and an eccentricity of 0.3 (da Silva et al., 2006; Pepper et al., 2020). In this paper, we first combined Rossiter-McLaughlin observations from the Keck Planet Finder (KPF) and stellar rotation periods derived from Transiting Exoplanet Survey Satellite (TESS) light curves to determine the true spin-orbit angle of HD 118203 b. Furthermore, we present the discovery of HD 118203 c, a long-period giant planet orbiting outside HD 118203 b. We characterized the three-dimensional orbital parameters of this outer planet by combining RVs and Hipparcos and Gaia astrometric data. HD 118203 is one of the first systems where the spin-orbit angle for both the hot Jupiter and the outer giant planet, and the mutual inclination between the hot Jupiter and outer planet have been determined. Recently, Maciejewski et al. (2024) presented an independent detection of HD 118203 c using long-baseline radial velocity data from HARPS-N (Cosentino et al., 2012) and the High-Resolution Spectrograph (HRS; Tull 1998) at the Hobby-Eberly Telescope (Ramsey et al., 1998) 222 Maciejewski et al. (2024), working in parallel with this work, also discovered HD 118203 c around the same time. The early draft of this work was completed and submitted before the findings of Maciejewski et al. (2024) were made public. However, during the peer-review process, we incorporated RV data from their paper as soon as it became available to more precisely characterize the system.. We compare our results with their findings in Section 5.3.
@@ 3.1 Literature values
HD 118203 is a sub-giant G0V star at an age of $5.32^{+0.96}_{-0.73}$ Gyr and a distance of $91.811\pm{0.236}$ pc (Pepper et al., 2020). The star has a mass of $1.25\pm{0.06}\ M_{\odot}$, a radius of $2.10\pm{0.06}\ R_{\odot}$, and a metallicity [Fe/H] of $0.223\pm{0.076}$ dex. Its spectroscopically determined projected rotation velocity, $v\sin i_{*}$, is $5.32\pm{0.5}$ km s-1 (Luck, 2017). Pepper et al. (2020) reported a rotation period of $20\pm{5}$ days from TESS light curves of sector 15 and 16. Recently, Castro-González et al. (2024) reported a period of 6.1 days from TESS PDCSAP light curves, which matches with the orbital period of HD 118203 b. They argue it could be resulted from the magnetic interaction between the hot Jupiter and the host star.
@@ 5.2 RV-only Keplerian Fit
We performed the Markov Chain Monte Carlo (MCMC) exploration with emcee (Foreman-Mackey et al., 2013) to estimate parameter credible levels. Our MCMC analysis used 50 walkers and ran for $1.5\times 10^{6}$ steps per walker, achieving a maximum Gelman-Rubin(GR) statistic of 1.004. Figure 6 shows the best fit Keplerian solution. We derived a minimum mass for HD 118203 c from RV amplitudes as $m_{c}\sin{c}=11.70^{+0.87}_{-0.83}\ \rm M_{J}$, semi-major axis as $a_{c}=6.32^{+0.17}_{-0.18}$ AU and an eccentricity of $e_{c}=0.268^{+0.028}_{-0.026}$. The derived planetary parameters are given in Table 5.
@@ 5.3 Three-dimensional orbit fitting
Semi-major axis $a_{c}$ (AU) $6.28^{+0.10}_{-0.11}$ no prior
@@
Figure 8 presents the joint posterior distributions for the orbital inclination, mass and orbital period of HD 118203 c (see Appendix A.4 for additional parameters). Our results show that HD 118203 c is a giant planet with a mass of $m_{c}=11.79^{+0.69}_{-0.63}\ \mathrm{M_{J}}$ semi-major axis of $a_{c}=6.28^{+0.10}_{-0.11}$ AU and eccentricity $e_{c}=0.26^{+0.03}_{-0.02}$. We obtained the orbital inclination of the outer giant planet as $I_{c}=93^{\circ}.7^{+11.9}_{-13.8}$, suggesting that HD 118203 c orbits the star in an almost edge-on orbit. Our results are consistent with the values reported from Maciejewski et al. (2024) within $1\sigma$.
Figure 9: Panel a: sketch showing the orbital geometry for the mutual inclination between HD 118203 b&\&&c. The observer is looking down from the z-axis. The x-y plane is the plane of the sky, and the star is centered on the origin. The orbital inclination of a planet is defined as the angle between the planet orbital axis and line of sight direction (z-axis). The mutual inclination is the angle between the orbital axes of the two planets. Panel b: For a transiting planet with orbital inclination $I_{1}=90^{\circ}$ without knowing the longitude of acceding node, its orbital axis could fall anywhere along the sky plane (orange). For a planet with an orbital inclination $I_{2}\neq 90^{\circ}$ without knowing the longitude of acceding node, its orbital axis falls on the surface of a cone (blue) with a vertex angle equals to the orbital inclination (toward observer, $I_{2}<90^{\circ}$) or $\pi-$ orbital inclination (away from observer, $I_{2}>90^{\circ}$).
Figure 10: (a) Orbital inclination of HD 118203 c ($I_{c}$, blue) , compared with that of HD 118203 b ($I_{b}$, red) and the line-of-sight stellar inclination ($i_{*}$, orange). Zero degrees correspond to the line of sight. The stellar inclination ranges from $0^{\circ}$ to $180^{\circ}$ to accounts for scenarios where the stellar axis may point either toward ($<90^{\circ}$) or away ($>90^{\circ}$) from the observer. (b): Distribution of HD 118203 b&\&&c mutual inclination projected in line of sight direction. (c): Distribution of spin-orbit angle of HD 118203 c projected in line of sight direction.
@@ 5.4 Constraints on the planet-planet mutual inclination and spin-orbit angle for HD 118203 c
where $I_{i}$ and $\Omega_{i}$ is the inclination and longitude of ascending node for planet $i$. Typically, using transit and radial velocity data, we can infer the inclination of a transiting planet, but not its longitude of the ascending node. In such scenarios, the orbital axis of a transiting planet with an inclination close to $90^{\circ}$ might be oriented in any direction within the sky plane. Similarly, for a non-transiting planet with an inclination away from $90^{\circ}$, its orbital axis could lie on the surface of a cone, which is either towards ($I_{2}<90^{\circ}$) or away from ($I_{2}>90^{\circ}$) the observer. Therefore, if we only have information about the orbital inclinations of two planets and the longitude of ascending nodes for one or both planets is unknown, we cannot determine the true mutual inclinations. Instead, we can only derive the mutual inclinations as projected along the line of sight.
@@ 6.2.1 Eccentricity excitation
However, our findings that the outer planet has an eccentricity of $e_{c}=0.26$, imply that we need to take into account the effects of the non-circular outer orbit. Naoz et al. (2011) explored a more generalized mechanism in which the outer orbit can be eccentric, called the eccentric Kozai-Lidov (EKL) mechanism. They found that considering an eccentric outer orbit requires an expansion of the perturbing function to octupole order, which is proportional to $(a_{1}/a_{2})^{3}$. In the octupole level of approximation, EKL mechanism can drive the eccentricity of the inner orbit to extremely high levels, and can even induce a complete reversal in the orientation of the inner orbit, flipping it from prograde to retrograde. The EKL mechanism is also applicable across a broader range of parameters, including scenarios with low initial mutual inclinations between the two objects. Li et al. (2014) demonstrated that when the two planets have eccentric orbits, the inner planet’s eccentricity can grow to large values even when the orbits are initially close to being coplanar. Petrovich (2015) investigated the consequence of this eccentricity excitation in the context of tides and found that subsequent high-eccentricity migration kept the two planets nearly coplanar. According to their study, hot Jupiters formed via this coplanar high-eccentricity migration (CHEM) are likely to have distant ( $a\gtrsim 5\rm{AU}$) and massive (three times more massive than the hot Jupiter) companions with relatively low mutual inclinations ( $\leq 20^{\circ}$) and moderate eccentricities (0.2-0.5). Our measurements for HD 118203 c, with properties of $a_{c}=6.28^{+0.10}_{-0.11}\ \rm{AU}$, $m_{c}=11.79^{+0.69}_{-0.63}\ M_{J}$, and $e_{c}=0.26^{+0.03}_{-0.02}$, align well with these predictions. Furthermore, our analysis indicates that $\approx 95\%$ values from the posterior distribution of the line-of-sight $\Delta I_{bc}$ are below $25^{\circ}$, consistent with the outer giant planet having a low mutual inclination relative to the inner hot Jupiter.
@@ 7 Conclusions
  • HD 118203 c is a long-period giant planet ($m_{c}=11.79^{+0.69}_{-0.63}\ \mathrm{M_{J}}$, $a_{c}=6.28^{+0.10}_{-0.11}$ AU, $e_{c}=0.26^{+0.03}_{-0.02}$) with moderate eccentricity outside the hot Jupiter. We constrained the orbital inclination of HD 118203 c ($I_{c}=93^{\circ}.7^{+11.9}_{-13.8}$ ) by combing RV and Hipparcos-Gaia astrometric accelerations. The line-of-sight mutual inclination between HD 118203 b &\&& c is $9^{\circ}.8^{+8.6}_{-6.7}$ at $1\sigma$ and $9^{\circ}.8^{+16.2}_{-9.3}$ at $2\sigma$ confidence level. The line-of-sight spin-orbit angle for HD 118203 c is $14^{\circ}.7^{+13.3}_{-10.2}$ at the $1\sigma$ confidence level and $14^{\circ}.7^{+26.4}_{-13.9}$ at the $2\sigma$ level confidence level. Our measurements demonstrate that the orbit of HD 118203 c is most likely aligned with both the host star spin axis and the orbit of the inner hot Jupiter.